# Nonuniform Sobolev Spaces

### 05 Jan 2024 v1

@@ 3.1. Embedding Theorem for Nonuniform Fractional Sobolev Spaces
$\int_{\mathbb{R}}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$ $\gtrsim\int_{a\geq 2}\frac{\mathrm{d}a}{a^{p_{1}(\alpha+s)-N+1}}=\infty.$
(ii)  Now assume that $p_{0}>p_{1}$. Take some $\varepsilon>0$ such that $\varepsilon<\min\{s,N/p_{1}-N/p_{0}\}$. Set $\alpha=N/p_{1}-\varepsilon$. We have $p_{1}\alpha<N<p_{0}\alpha$ and $p_{1}(\alpha+s)=N+p_{1}(s-\varepsilon)>N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{s}^{p_{1}}$.
(iii)  When $\tilde{s}<N/p_{1}-N/p_{0}$, we have $p_{0}>p_{1}$. Take some constant $\alpha$ such that
$\max\Big{\{}\frac{N}{p_{0}},\frac{N}{p_{1}}-s\Big{\}}<\alpha<\frac{N}{p_{1}}-\tilde{s}.$
Then $p_{0}\alpha>N>p_{1}\alpha$, $p_{1}(\alpha+s)>N$ and $p_{1}(\alpha+\tilde{s})<N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{\tilde{s}}^{\vec{p}}$. ∎
@@ 3.2. Density of Compactly Supported Infinitely Differentiable Functions
we derive from Minkowski’s inequality and the continuity of translation operators in $L^{p_{0}}$ that
$\lim_{\lambda\rightarrow 0}\|g_{\lambda}-f\|_{L^{p_{0}}}=0$ (3.19)
@@
$[g_{\lambda}]_{W_{s}^{p_{1}}}$ $=\bigg{\|}\frac{g_{\lambda}(x)-g_{\lambda}(z)}{|x-z|^{N/p_{1}+s}}\bigg{\|}_{L^{p_{1}}(\mathbb{R}^{N}\times\mathbb{R}^{N})}\leq\int_{\mathbb{R}^{N}}[f]_{W_{s}^{p_{1}}}\varphi(y)\mathrm{d}y=[f]_{W_{s}^{p_{1}}},\quad\forall\lambda>0.$
On the other hand, we see from (3.19) that there is a subsequence $\{g_{\lambda_{l}}:\,l\geq 1\}$ which is convergent to $f$ almost everywhere. It follows from Fatou’s lemma that
$[f]_{W_{s}^{p_{1}}}\leq\liminf_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}\leq\limsup_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}\leq[f]_{W_{s}^{p_{1}}}.$
Hence
$\lim_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}=[f]_{W_{s}^{p_{1}}}.$
That is, $\{(g_{\lambda_{l}}(x)-g_{\lambda_{l}}(z))/|x-z|^{N/p_{1}+s}:\,l\geq 1\}$ converges to $(f(x)-f(z))/|x-z|^{N/p_{1}+s}$ both in $L^{p_{1}}$ norm and almost everywhere on $\mathbb{R}^{N}\times\mathbb{R}^{N}$. It follows from Riesz’s theorem that
$\lim_{l\rightarrow\infty}[g_{\lambda_{l}}-f]_{W_{s}^{\vec{p}}}=0.$
Together with (3.19), we get the conclusion as desired. ∎

### 23 Jan 2024 v3

@@ 3.1. Embedding Theorem for Nonuniform Fractional Sobolev Spaces
$\int_{\mathbb{R}}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$ $\gtrsim\int_{a\geq 2}\frac{\mathrm{d}a}{a^{p_{1}(\delta+s)-N+1}}=\infty.$
(ii)  Now assume that $p_{0}>p_{1}$. Take some $\varepsilon>0$ such that $\varepsilon<\min\{s,N/p_{1}-N/p_{0}\}$. Set $\delta=N/p_{1}-\varepsilon$. We have $p_{1}\delta<N<p_{0}\delta$ and $p_{1}(\delta+s)=N+p_{1}(s-\varepsilon)>N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{s}^{p_{1}}$.
(iii)  When $\tilde{s}<N/p_{1}-N/p_{0}$, we have $p_{0}>p_{1}$. Take some constant $\delta$ such that
$\max\Big{\{}\frac{N}{p_{0}},\frac{N}{p_{1}}-s\Big{\}}<\delta<\frac{N}{p_{1}}-\tilde{s}.$
Then $p_{0}\delta>N>p_{1}\delta$, $p_{1}(\delta+s)>N$ and $p_{1}(\delta+\tilde{s})<N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{\tilde{s}}^{\vec{p}}$. ∎
@@ 3.2. Density of Compactly Supported Infinitely Differentiable Functions
we have
$g_{\lambda}(x)-f(x)=\int_{\mathbb{R}^{N}}(f(x-\lambda y)-f(x))\varphi(y)\mathrm{d}y.$
Now we derive from Minkowski’s inequality and the continuity of translation operators in Lebesgue spaces that
$\lim_{\lambda\rightarrow 0}\|g_{\lambda}-f\|_{L^{p_{0}}}\leq\lim_{\lambda\rightarrow 0}\int_{\mathbb{R}^{N}}\|f(\cdot-\lambda y)-f\|_{L^{p_{0}}}\varphi(y)\mathrm{d}y=0$ (3.19)
@@
$\lim_{\lambda\rightarrow 0}[g_{\lambda}-f]_{W_{s}^{p_{1}}}$
$\leq\lim_{\lambda\rightarrow 0}\int_{\mathbb{R}^{N}}\bigg{\|}\frac{\big{(}f(x-\lambda y)-f(x)\big{)}-\big{(}f(z-\lambda y)-f(z)\big{)}}{|x-z|^{N/p_{1}+s}}\bigg{\|}_{L_{(x,z)}^{p_{1}}(\mathbb{R}^{N}\times\mathbb{R}^{N})}\varphi(y)\mathrm{d}y$
$=\lim_{\lambda\rightarrow 0}\int_{\mathbb{R}^{N}}\bigg{\|}\frac{f(x-\lambda y)-f(z-\lambda y)}{|(x-\lambda y)-(z-\lambda y)|^{N/p_{1}+s}}-\frac{f(x)-f(z)}{|x-z|^{N/p_{1}+s}}\bigg{\|}_{L_{(x,z)}^{p_{1}}(\mathbb{R}^{N}\times\mathbb{R}^{N})}\varphi(y)\mathrm{d}y$
$=0,$
having applied the fact that $(f(x)-f(z))/|x-z|^{N/p_{1}+s}\in L_{(x,z)}^{p_{1}}$. Hence $\lim_{\lambda\rightarrow 0}\|g_{\lambda}-f\|_{W_{s}^{\vec{p}}}=0$. ∎